OpenStudy (anonymous):

HELP PLEASE!!!! A spherical storage tank has a diameter of 14 feet. Approximately how many cubic feet of water will it hold?

OpenStudy (mathstudent55):

The volume of a sphere with r radius is: \( \large{ V = \dfrac{4}{3}\pi r^3 } \) You are given the diameter of the sphere, The radius is half the diameter: \( \large { radius = r = \dfrac{diameter}{2} }\)

OpenStudy (jasmineflvs):

1436.76 cubic feet would be its volume

OpenStudy (jasmineflvs):

is that right??

OpenStudy (mathstudent55):

Yes

OpenStudy (jasmineflvs):

thanks

OpenStudy (mathstudent55):

yw